OpenStudy (anonymous):

Solve -5(3x-12) =30 - 9x please show all your work

OpenStudy (callisto):

-5(3x-12) =30 - 9x -15x + 60 = 30 -9x 6x = 30 can you solve it then?

OpenStudy (anonymous):

?????

OpenStudy (callisto):

solve the answer yourself, only 1 step is left for you. at least you should try?